OpenStudy (anonymous):

I need to determine the rank of this matrix. I have tried it, but I get lost with the row operations. Any tips and trick how to do it?

OpenStudy (anonymous):

I need to determine the rank of the matrix for every \[(a,b) \epsilon \mathbb{R} ^{2}\]

OpenStudy (rational):

first step : divide the 3rd row by (a+1)

OpenStudy (rational):

second step : R2 - R1

OpenStudy (loser66):

\(\left[\begin{matrix}1&(a-1)&2&(a+2)&(a+b)\\1&2a&0&a&2a+b\\0&-(a+1)&(2a+2)&0&0\\0&(2a+2)&(4a-4)&(a^2+a-8)&(4a+ab+b)\end{matrix}\right]\) \(R_3*2 +R_4-->R_4\) \(\left[\begin{matrix}1&(a-1)&2&(a+2)&(a+b)\\1&2a&0&a&2a+b\\0&-(a+1)&(2a+2)&0&0\\0&0&((4a)^2-16)&(a^2+a-8)&(4a+ab+b)\end{matrix}\right]\) \(R_2/(a+1)\)--> \(\left[\begin{matrix}1&(a-1)&2&(a+2)&(a+b)\\1&2a&0&a&2a+b\\0&-1&2&0&0\\0&0&((4a)^2-16)&(a^2+a-8)&(4a+ab+b)\end{matrix}\right]\) \(R_1*(-1)+R_2)\)--> \(\left[\begin{matrix}1&(a-1)&2&(a+2)&(a+b)\\0&(1+a)&-2&-2&a\\0&-1&2&0&0\\0&0&((4a)^2-16)&(a^2+a-8)&(4a+ab+b)\end{matrix}\right]\)

OpenStudy (loser66):

\(R_3+R_2-->R_2\)--> \(\left[\begin{matrix}1&(a-1)&2&(a+2)&(a+b)\\0&a&0&-2&a\\0&-1&2&0&0\\0&0&((4a)^2-16)&(a^2+a-8)&(4a+ab+b)\end{matrix}\right]\) \((R_2 +R_3)(-1)+R_1-->R_1\) \(\left[\begin{matrix}1&0&0&(a+4)&b\\0&a&0&-2&a\\0&-1&2&0&0\\0&0&((4a)^2-16)&(a^2+a-8)&(4a+ab+b)\end{matrix}\right]\)

OpenStudy (loser66):

That's all I can do for you. They are linearly independent, so that rank A =4

OpenStudy (anonymous):

Thanks @Loser66 - I'll try to go through the steps now!